OpenStudy (anonymous):

Find the domain : f (x) = 3x - 8 divide 4x + 20

OpenStudy (anonymous):

The domain is all the values that x can be in the equation. If 4x + 20 is the whole denominator of this function, then the only place where we would have trouble is where 4x + 20 = 0 or when x = -5. If you chose x = -5, you would be dividing by zero which we don't want. So, the domain works for all real numbers except for -5. In interval notation, that's: \[(-\infty,-5)\cup(-5,\infty)\]

OpenStudy (anonymous):

ty